A single stock futures contract on a non-dividend paying stock with current price $110 has a maturity of one year.

a. If the T-bill rate is 4.0%, what should the futures price be? (Round your answer to 2 decimal places.)

b. What should the futures price be if the T-bill rate is still 4.0% and the maturity of the contract is three years? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

c. What if the interest rate is 5.1% and the maturity of the contract is three years? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Futures price=Stock price*(1+interest rate)^t

1.
=110*(1+4%)=114.4

2.
=110*(1+4%)^3=123.73504

3.
=110*(1+5.1%)^3=127.70292161